This webpage contains information on the maintenance procedures of the optics systems of the ING.
Optical action list
| Date when added | Action | Fault number | Status | Comment |
| 15/11/2006 | ADC has to be separated, and glued together again | pending | The two prisms are not identical, there was deviation of the laser beam, when it was measured in the optics lab at SLO. Probably the gluing cause the problem. The manufacturer was ICOS. |
| Date when added | Action | Fault number | Status | Comment | Date when finished |
| 22/02/2007 | OASIS 33mm enlarger | Done | The OASIS 33mm enlarger problem had been investigated. With the removal of the mask it was possible to eliminate vignetting and so the observations with the 33mm are now possible. Due to the removal of the mask some scattered light was detected. At the moment the mask is out of OASIS. For certain observations (faint targets) it could be important to reduce the scattered light, in this case well in advance (2-3 month) the T&I group should be contacted to do the job. Although due to the difficulty of the mask removal and installation is not recommended. | 2010 | |
|
01/04/2007
|
Scattered light in NAOMI
|
Done
|
The source of
the problem was found, the OSCA electronics module has
some LED, which caused the light contamination in GRACE.
Some work has been done to eliminate the scattered light previously:
|
2009 | |
| 03/01/2008 | JKT primary and secondary mirror vashing | Done | Reflectivity_results_page | 03/01/2008 | |
| 13/12/2007 | INT primary mirror vapour washing | Done | Reflectivity_results_page | 13/12/2007 | |
| 05/11/2007 | ISIS red fold mirror change | Done | During the ISIS alignment checks it turned out that the red fold mirror was in a terrible state. We decided to change it. Two fold mirrors were found at sea-level. The new red fold mirror was installed successfully. There is one more spare red fold mirror, it is stored in the General Optics cabinet (the floor under the Observing floor). | 10/11/2007 | |
| 05/11/2007 | ISIS alignment checks | Done | ISIS was brought to the test focal station and the alignment of the optical components were checked based on the alignment documents. The tools for the alignment were put back to their cabinets labelled "ISIS alignment tools" on the mezzaine floor (under the Observing floor). Some of the tools were so big they are stored close to the cabinets, but only on the floor or on the top of other cabinets. | 23/11/2007 | |
| 29/10/2007 | WHT visual check | Done | Reflectivity_results_page | 29/10/2007 | |
| 04/10/2007 | WHT primary mirror snowcleaning | Done | Reflectivity_results_page | 04/10/2007 | |
| 04/10/2007 | INT primary mirror snowcleaning | Done | Reflectivity_results_page | 04/10/2007 | |
| 03/10/2007 | Problem regarding screw and washers for the filter adaptors for the PFIP filter wheel (see comments). | #17861 | Done |
One of the circular 76mm
Taurus filter adaptors was missing screws and its
retaining ring. 2 of the 50mm square filter adaptors
were missing screws and washers. These adaptors are
located in the WHT Prime Cabinet.
The 50mm square filter adaptors have now the screws and plastic washers, so they are fine. Regarding the 76mm Taurus filter holders, the retaining ring is still missing, but the screws and the washers are in their place. The washers are taped, so they can hold the filters inside the adaptor, so it also fine and could be used. |
03/10/2007 |
| 29/09/2007 | The filter coating on the Stromgen B filter (#200) for the WFC is scratched. | #17854 | Investigate more |
The filter coating on the
Stromgen B filter (#200) for the WFC is seriously
scratched away over a large portion of the filter area. After the inspection of the filter it could be seen that it is indeed seriously scratched and probably this affect the performance of the filter. The reorder of the filter is really difficult and extremely expensive (low number filters to be ordered - high cost) according to the experiences. Since we can not have a new filter, the only thing that could be done is the measurement of the filter`s spectral characteristics and comparison to previous data. |
01/10/2007 |
| 20/09/2007 | Problem with image quality in U filter at WFC at INT. | #17840 | investigate more | Problem with image quality in U filter at WFC at INT. I prepared (Tibor) a document about INT image quality within 40 x 40 arcmin for 360nm and 650nm. I analyzed the nominal configuration and one, in which the corrector lenses were shifted. It could be seen that the spots are increased significantly above 15 arcmin field radius in this latter case for both wavelengths. The spot diagrams for the nominal configuration were fine throughout the whole 20arcmin field radius (40` x 40`). According to the Zemax file the primary was tilted with 0.0001degree and the PF corrector lenses were shifted along the optical axis with 3mm in 2005.07.19. It seems that the spot elongation is the effect of the distortions introduced by the shift of the corrector lenses. (The tilt of the primary is not a problem, if we only take into account the 0.0001degree mirror tilt then the diagrams are the same with the original.) Questions: Do we know something about this 3mm shift of the 3 corrector lenses? Why was it applied exactly? How this image degradation above 15 arcmin radius affect the observations? According to the analysis the image quality degraded in the 650nm case as well. | 22/11/2007 |
| 15/09/2007 | The filter nr 198 in INT WFC is loose in its holder. | #17824 | Done | The filter nr 198 in INT WFC is loose in its holder. The filter is fixed in its position, it could be used now. | 29/10/2007 |
| 27/07/2007 | Order 2pcs of GG495 | Done | During cleaning of the filters it turned out that there are scratches on the filter. It was necessary to order new filters; 2 pcs from GG495 (Size: 80mm x 35mm x 3mm) | 24/09/2007 | |
| 27/07/2007 | ISIS GG495 filter is dusty (Order sorting filter for ISIS) | #17774 | Done | It is cleaned, but there are scratches on the filter. It is necessary to order new filters; at least 2 pcs from GG495 | 27/07/2007 |
| 24/07/2007 | WHT primary mirror snowcleaning | Done | Reflectivity_results_page | 24/07/2007 | |
| 24/07/2007 | INT primary mirror snowcleaning | Done | Reflectivity_results_page | 24/07/2007 | |
| 19/06/2007 | In ISIS Red Filter A (RFA) the GG495 filter is not mounted properly | Done | Mechanical work is needed to correct this. It is solved. | 20/07/2007 | |
| 30/04/2007 | Order 2 pcs of new laser goggles | Done |
Send back the one wrong
goggles, exchange it to the Art No.018.T0027.00, with
the reinforced protector frame; order one more glass
with Art No. 018.T0027.00 filter and reinforced
protector frame Company: http://www.lvg.com/ Contact: Tim Frieb |
29/06/2007 | |
| 13/02/2007 | WYFFOS scattering analysis | Done |
There are two different
places were identified, where if proper baffles are
installed, the scattered light could be reduced. The
analysis showed that apart from the scattered light,
ghost images are also present. These could not be
eliminated with baffling.
The report about the tests are here: |
07/06/2007 | |
| 22/06/2007 | INT primary mirror snowcleaning | Done | Reflectivity_results_page | 22/06/2007 | |
| 26/05/2007 | WYFFOS relay mirror installation and alignment | Done | The freshly recoated mirror is installed and aligned in WYFFOS | 31/05/2007 | |
| 13/12/2006 | WYFFOS relay mirror has arrived (it was sent to Denton US because of recoating) | Done | Denton US, blue enhanced silver coating, DENTON FSS99-500 coating | 17/05/2007 | |
| 02/04/2007 | Install the new filters at INT (IDS/A&G ASCF filters) | Done |
The current positions and
filters: position, filter type 0, Clear 1, GG385 2, GG395 3, GG495 4, RG630 5, BG28 (We have one more type of filter - BG38 -, which is not in the filter holder) Filters: http://whtics.roque.ing.iac.es:8081/FilterDatabase/ |
02/05/2007 | |
| 13/04/2007 | WHT primary mirror snowcleaning | Done | Reflectivity_results_page | 13/04/2007 | |
| 09/04/2007 | Laser goggles investigation, labelling | Done |
Those laser goggles that
could be used for GLAS has to be separated from the
others. All of the laser goggles have to be labelled. Specifications: Optical documents webpage |
11/04/2007 | |
| 01/02/2007 | Order First Contact Deluxe Kit | Done |
It was ordered from LASER
2000 (www.laser2000.es) Contact person: Berenice ALANIS-PIGEYRE (alanis@laser2000.fr) |
28/02/2007 | |
| 24/01/2007 | Test the First Contact cleaning liquid on different optical components | Done |
http://www.photoniccleaning.com/ Tests were carried out on different optical components. The result that it could be used on smaller optical surfaces, eg on mirrors, filters, smaller lenses. It contains acetone, which is highly flammable, so it is not suggested to clean with this method on larger surfaces (though the company states that it is possible). Also not recommended to use on gratings. |
01/02/2007 | |
| 22/01/2007 |
Order of the following
filters (2 pcs of each): GG385, GG395, GG495, BG38, BG28, RG630. Their size is: 60x45x2mm |
Done |
The items were ordered
from ITOS (www.itos.de)
Contact person: Sabine Sollinger (sollinger@itos.de) Other company for ordering filters: Linos (www.linos.com) |
02/04/2007 | |
|
16/01/2007
|
Quotations on reflectivity, scattering and surface roughness measuring devices are needed. Also it is needed to contact SMS to get info about the reflectivity meter's not reliable operation. | Done |
Companies offering
reflectivity meters: 1. Schmitt (http://www.schmitt-ind.com/products-services-measurement-systems.shtml) 2. Konica Minolta (http://www.konicaminolta.eu/products/industrial_products/overview) 3. DMO (http://www.dmo.be/web/framehtm.htm) |
25/01/2007 | |
| 29/01/2007 | Aluminising the primary mirror of the WHT | Done | Reflectivity_results_page | 29/01/2007 | |
| 18/01/2007 | INT primary mirror snowcleaning | Done | Reflectivity_results_page | 18/01/2007 | |
| 18/01/2007 | Isis slit and dekker cleaning | #17513 | Done |
The Isis slit and
dekker was cleaned with Balzers 2 liquid, optical tissue
and gloves were used during the cleaning process The optical cleaning elements could be found in the yellow cabinet next to the sink at the basement of WHT |
18/01/2007 |
| 08/01/2007 | WHT primary snowcleaning | Done | Reflectivity_results_page | 08/01/2007 | |
| 20/12/2006 | WHT primary mirror reflectivity measurements | Done | Reflectivity_results_page | 20/12/2006 | |
| 05/12/2006 | WHT primary mirror reflectivity measurements | Done | Reflectivity_results_page | 05/12/2006 | |
| 28/11/2006 | Vapour washing of the WHT primary mirror with a new technique | Done |
The technique was tested,
one day we washed only the half of the mirror (on the
28th - right side, 29th - left side, 30th - right side
again as seen from AP3). Results could be found here:
And information on the process here: |
28-29-30/11/2006 | |
| 02/11/2006 | WYFFOS spherical mirror reflectivity measurement | Done | Results could be found here: | 02/11/2006 | |
| 02/11/2006 | WYFFOS relay mirror change - use the spare | Done | The WYFFOS relay mirror has to be sent for recoating (Denton US, blue enhanced silver coating, DENTON FSS99-500 coating needed) | 02/11/2006 | |
| 17/02/2003 | CO2 cleaning WHT, INT, JKT | ||||
| 03/02/2003 | Reflectivity measurement WHT, INT, JKT | ||||
| 20/01/2003 | CO2 cleaning WHT, INT, JKT | ||||
| 06/01/2003 | Reflectivity measurement WHT, INT, JKT | ||||
| 23/12/2002 | CO2 cleaning INT, JKT | ||||
| 09/12/2002 | Reflectivity measurement WHT, INT, JKT | ||||
| 16/09/2002 | Filter measurement | Done |
Measure the
TAURUS filters to be used in the WHT PFC History: 17/09/2002 Checked the spectrograph at SLO.It is not working correctly. Varian is contacted, waiting for reply. 20/09/2002 contacted Varian for support. 20/11/2002 Ordered new lamp from Varian.
|
||
| 12/06/2002 | Update Filter database | Done |
History: 12/06/2002 Did not update filter database. Waiting till new version is out. Contacts Robert Greimel and Craige Bevil. |
||
| 12/06/2002 | Filter check | Done |
History: 12/06/2002 Checked filters and cleaned when necesary. |
||
| 15-17/12/2002 | WHT aluminising, Nasmyth washing | Done | Reflectivity_results_page | 15-17/12/2002 | |
| 23/11/2002 | WHT, INT, JKT CO2 cleaning | Done |
Reflectivity_results_page Could not take pictures of WHT because of GRACE building |
28/11/2002 | |
| 04/11/2002 | Done | Reflectivity_results_page | 14/11/2002 | ||
| 28/10/2002 | WHT, INT CO2 cleaning | Done | Reflectivity_results_page | 30/10/2002 | |
| 24/10/2002 | JKT washing | Done | Reflectivity_results_page | 24/10/2002 | |
| 07/10/2002 | Reflectivity measurement WHT, INT, JKT | Done | Reflectivity_results_page | 08/10/2002 | |
| 02/10/2002 | INGRID alignment check with Sue Worswick | PIMS issue 121 | Found a detector tilt of 400 microns over the y direction and 100 microns over the x directrion. Also have to look at the PS positions. | 08/10/2002 | |
| 23/09/2002 | CO2 cleaning WHT, INT, JKT | Reflectivity_results_page | 25/09/2002 | ||
| 16/09/2002 | Optical stock WHT cabinet | Done | Need optical tissue is WHT cabinet | 25/09/2002 | |
| 13/09/2002 | Filter check INT WFC | #15304 | Done |
13/09/2002
Ask Danny Lennon if orientation fix does not change
science output. History: 13/09/2002 checked filter orientation. A few filters are wrongly orientated.25/09/2002 Re orientated filters #199, #207, #210, #220
|
25/09/2002 |
|
13/09/2002 |
Reflectivity measurement WHT, INT, JKT | Reflectivity_results_page | 13/09/2002 | ||
| 13/08/2002 | Clean corrector plate of RoboDIMM. | Done washed it with water | 25/09/2002 |